OpenStudy (anonymous):

what is the difference of these mixed numbers ? 8 2/5 - 6 1/10 I believe it might be 2 2/5 idk

OpenStudy (anonymous):

2 and 3/10

OpenStudy (anonymous):

2 3/10

OpenStudy (turingtest):

8 2/5=42/5 6 1/10=61/10 42/5-61/10=(84-61)/10=23/10=2 3/10

OpenStudy (anonymous):

yes each step counts thank u